undeadknight26 (undeadknight26):

Meg enjoys listening to classical music and downloads music files regularly from the internet. She already has three hours of classical music in her collection. From the second month onwards, the number of hours of classical music she downloads each month is five percent greater than the number of hours of classical music she downloaded the previous month. Which function best shows the total number of hours of classical music Meg has in her collection f(x) after x months? f(x) = 2(1.05)^x f(x) = (1.05)^x + 3 f(x) = (1.05)^x - 3 f(x) = (1.05)^x + 2

undeadknight26 (undeadknight26):

@justicehunter i think it is the second one.

OpenStudy (anonymous):

yes thats correct

undeadknight26 (undeadknight26):

The function f(x) is shown below. f(x) = (3)x - 2 The function f(x) is shifted to the left by 13 units. Which of the following best represents the new function? f(x) = (3)^x + 11 f(x) = (3)^x + 13 f(x) = (3)^x -15 f(x) = (3)^x - 13

undeadknight26 (undeadknight26):

@justicehunter

undeadknight26 (undeadknight26):

i think it was the last one.

undeadknight26 (undeadknight26):

@hartnn @E.ali @AllTehMaffs @abb0t @charlotte123 @CGGURUMANJUNATH please help

OpenStudy (anonymous):

and f(x) = (3)x - 2 is this f(x)=3^(x-2) ?

undeadknight26 (undeadknight26):

oh i thought it was to subtract to shift it left and add to shift it right

OpenStudy (anonymous):

for example y=x. if you write y=x+13 then you shift the graph 13 units left, correct?

undeadknight26 (undeadknight26):

so it was the second one?

OpenStudy (anonymous):

no you can always try how things work on y=x its simple you can graph y=x+13 and see that the graph is shifted 13 units left. that indicates that f(x+13) is the answer!

OpenStudy (anonymous):

f(x)=3^(x-2) ? is f(x) this? you wrote it wrong on the question

undeadknight26 (undeadknight26):

so it has to be the second one then because these are the only choices?

undeadknight26 (undeadknight26):

there are no parenthises i copy and pasted..

OpenStudy (anonymous):

no but there is no exponent on the question you see.

OpenStudy (anonymous):

f(x) = (3)x - 2 this mean 3x-2 you always have to use correct notation so we can understand the question:(

undeadknight26 (undeadknight26):

lol this is all it says...

OpenStudy (anonymous):

if x-2 is exponent you have to write 3^(x-2) it like that. so instead of x put x+13 and there is your answer

undeadknight26 (undeadknight26):

it was F(x) = (3)^x - 2

OpenStudy (anonymous):

yes when you put x+13 instead of x you'll end up with 3^x+11

undeadknight26 (undeadknight26):

so which one is it???

OpenStudy (anonymous):

f(x) = (3)^x + 11 ? the first one

undeadknight26 (undeadknight26):

oh...lol...yea i didn't see that post up there...

undeadknight26 (undeadknight26):

ok thank you i have to submit the test now...the epic moment...

undeadknight26 (undeadknight26):

43/50 not bad thanks!

OpenStudy (anonymous):

no problem! i wish you the best!